Find the first seven values (i.e., x(n) for n = 0 to 6) of the function x(n)=[-nu(n)]+[2(n-3)u(n-3)]+[4u(n-4)]
Answer #1

given function, x Ch) - - un) + QT--3)+ + u6 - ) Til ... un) =fı nzo nzo - nao S-2 0 1 2 3 *40-3) oli n3 vm-l) - Si n23 - Toforn=Y XCt) = -4 (170) X4) = -4 4C4) + 24-3) 14-3) + 4 4(1-4) (170) 1-4X1 +2Q) 40 + 4460) in sonao =-4 + 2x! +4x1 = 21 for n=Please give me rating.